soybean meal is 14% protein, cornmeal is 7% protein, how many pounds of each should be mixed together in order to get a 280 pound mixture that is 12% protein
Answer by checkley77(12844)

.14x+.07(280-x)=.12*280
.14x+19.6-.07x=33.6
.07x=33.6-19.6
.07x=14
x=14/.07
x=200 pds. of soybeans is used.
280-200=80 pds. of cornmeal is uded.
Proof:
.14*200+.07*80=33.6
28+5.6=33.6
33.6=33.6